# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ Abstract.
Robotic Process Automation (RPA) offers a valuable solution for efficiently automating tasks on the graphical user interface (GUI), by emulating human interactions, without modifying existing code. However, its broader adoption is constrained by the need for expertise in both scripting languages and workflow design. To address this challenge, we present PromptRPA, a system designed to comprehend various task-related textual prompts (e.g., goals, procedures), thereby generating and performing corresponding RPA tasks. PromptRPA incorporates a suite of intelligent agents that mimic human cognitive functions, specializing in interpreting user intent, managing external information for RPA generation, and executing operations on smartphones. The agents can learn from user feedback and continuously improve their performance based on the accumulated knowledge. Experimental results indicated a performance jump from a 22.28% success rate in the baseline to 95.21% with PromptRPA, requiring an average of 1.66 user interventions for each new task. PromptRPA presents promising applications in fields such as tutorial creation, smart assistance, and customer service.
Robotic Process Automation, UI navigation, natural language, UI understanding
copyright: noneccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Interactive systems and tools
Figure 1. The workflow of PromptRPA. Blue arrows guide the generation process from textual prompts to RPA, while green arrows indicate the knowledge accumulation that occurs through user interactions, thereby enhancing the future performance of the agents.
@@ 1. Introduction
Robotic Process Automation (RPA) has emerged as a pivotal technology, significantly boosting productivity and operational efficiency by automating a wide range of tasks across various applications (Chakraborti et al., 2020). Its ability to simulate human interactions on the graphical user interface (GUI) allows for broad compatibility with third-party applications, without necessitating modifications to existing code (Jovanović et al., 2018). However, RPA primarily relies on predefined operation sequences (Ribeiro et al., 2021; Herm et al., 2023). Customizing these sequences often requires specialized skills in scripting languages and workflow design, which is a significant barrier to adoption, particularly for users without a technical background (Aguirre and Rodriguez, 2017). Additionally, the fixed operation sequences have difficulty adapting to the frequently updated GUI, making them susceptible to becoming quickly outdated.
@@
To address this gap, we introduce PromptRPA, which employs a text-prompt-driven RPA generation approach implemented on smartphones. PromptRPA is capable of accepting unrestricted textual descriptions for automation tasks—from concrete operational instructions like “open settings, locate WLAN, and…” to high-level functional requests such as “change the ringtone”. PromptRPA then automatically translates these textual prompts into executable operation sequences on GUI, freeing users from the complexity of designing automation scripts themselves.
@@
PromptRPA serves as an intelligent navigation system on smartphones. It interprets user intent from textual descriptions and matches them to its internal “map” for route planning and decision-making, thereby automatically generating RPA workflows. Should errors or issues arise, PromptRPA is designed to adapt dynamically to ensure uninterrupted operation. Throughout the process, PromptRPA faces several challenges. It needs to accommodate diverse and partially articulated textual descriptions, which requires inferring intentions from vague or incomplete instructions (Wen et al., 2001). Additionally, it is challenging to deal with its incomplete “map”, which may lack detailed route information for less common tasks or newly added application functions (Islam et al., 2010). Another challenge is adapting to frequent updates in interfaces and applications; as they evolve, the routes PromptRPA relies on may also require updates (Venkatesh et al., 2022). These challenges span across different stages of the process and require PromptRPA to be highly flexible and adaptable.
@@
To meet these challenges, we employ a multi-agent framework (Panait and Luke, 2005) for PromptRPA. The framework allocates specialized agents to sub-tasks within different stages of the process, with each agent emulating specific human cognitive functions. As illustrated in Figure 1, PromptRPA spans three stages—from information collection and instruction generation to operation mapping. Specifically, information collection involves gathering all necessary descriptions of functions and steps for the RPA task. Instruction generation entails creating formalized instructions to guide operations, and operation mapping refers to translating text-based instructions into operations on smartphones to ultimately complete the RPA task.
@@
To support the multi-agent framework, two core elements are critical: a comprehensive knowledge base (Venkatesh et al., 2022), which is our smartphone “map”, and stable decision-making (Zhong et al., 2021), which ensures the agents’ efficient and collaborative task execution. Therefore, we create an evolving knowledge base that expands as it accumulates data and user interactions. We introduce three key agents to meet the specific knowledge demands of RPA generation: the Analysis Agent, which actively fills in missing contextual parameters from user inputs; the Retrieval Agent, which sources online tutorials to guide task automation; and the Mobile Semantics Agent, which conducts in-depth semantic analysis of mobile GUI, for better alignment with textual inputs. To bolster decision-making, the Assessment Agent scrutinizes decisions made by other agents, serving as an internal audit mechanism to improve accuracy and reliability. Beyond internal audit, our agents are seamlessly integrated into a user interface, allowing users to oversee and enhance decisions. User usage and intervention create a feedback loop for continuous knowledge accumulation and system improvement, thereby optimizing efficiency for future tasks.
@@
To evaluate PromptRPA, we conducted a pipeline evaluation with 2,500 textual prompts crowd-sourced. The experiment demonstrated that PromptRPA increased the task success rate from a baseline of 22.28% to 95.21%, requiring only an average of 1.66 user interventions for each new task. Moreover, it was proved that the accumulation of user usage and intervention improved system performance. Furthermore, a user study involving both skilled and unskilled smartphone users confirmed PromptRPA’s effectiveness in assistance and its low interaction cost. These results substantiate PromptRPA’s feasibility and usability in practical settings.
Specifically, we propose PromptRPA, a multi-agent system accepting unrestricted textual expression of task requirements, supplementing with external knowledge primarily from online tutorials, and thus able to complete a wider range of RPA tasks. By presenting the results of each stage, PromptRPA supports more intuitive user interactions including dialogues, edits, and demonstrations. It effectively reduces user burden while ensuring a high success rate of RPA tasks. Due to the low-cost translation from text to RPA, PromptRPA has promising applications in generating interactive tutorials, enhancing smart assistance, and improving online customer service.
@@ 2. Related Work
Our work intersects with four research areas: prior RPA systems, the knowledge required for RPA, LLM-driven multi-agent frameworks, and optimizing the system based on user feedback.
2.1. Prior RPA Systems
Prior RPA systems on smartphones typically work by replaying predefined interaction flows on graphical user interfaces, thereby completing the corresponding tasks. Early systems like SUGILITE (Li et al., 2017) and APPINITE (Li et al., 2018) utilized programming by demonstration to enable users to custom automation flows. However, for complex or unfamiliar functions, creating flows is challenging, especially for beginners and the elderly, who may lack essential task execution skills (Chen et al., 2017; Jin et al., 2022). Subsequent developments introduce interactive, context-aware tutorials for GUI applications (Fraser et al., 2019; Pongnumkul et al., 2011; Li et al., 2015; Zhong et al., 2021; Li, 2021). However, these systems primarily rely on pre-configured flows by developers (Wang et al., 2014), and accommodating new tasks requires users to have specialized skills in script design (Wewerka et al., 2020). Moreover, these fixed flows become outdated when interface updates (Wu et al., 2023).
Recent efforts have focused on translating natural language into operation sequences on smartphones. Notable contributions include Branavan et al. (Branavan et al., 2009) and Li et al. (Li et al., 2020a). However, they require detailed instructions for each step, necessitating that users have a thorough understanding of task execution. While there are systems that support general task requests, limitations exist, such as UGIF’s reliance on a pre-collected tutorial dataset (Venkatesh et al., 2022) and Voicify’s dependence on app developers’ implementation of deep links (Vu et al., 2023), failing to meet a wide range of user needs. Additionally, they often function as a single “black box”, making intervention or correction difficult for users when errors occur (Venkatesh et al., 2022).
For easier use, our system is designed to accept unrestricted text inputs related to the task, without requiring users to have prior knowledge of task execution. By integrating online tutorial data and enhancing real-time analysis of the mobile context, our system more effectively adapts to diverse user needs and evolving interfaces. Moreover, our system features a user-centric, transparent design pipeline to facilitate low-effort user intervention.
2.2. Knowledge Required for RPA
@@
Addressing user input presents challenges in accuracy and scope. User inputs are often vague (Jansen et al., 1998), requiring supplementary contextual information for accurate recognition of user intent (Smith et al., 2006; Grudin and Jacques, 2019). Inspired by these studies, our system necessitates a deeper exploration of context information related to mobile usage. In terms of scope, existing UI trace datasets such as Rico (Deka et al., 2017), UGIF-DataSet (Venkatesh et al., 2022), and PixelHelp (Li et al., 2020a) are pre-collected and thus failing to meet a wide array of dynamic user needs (Beneteau et al., 2019). AutoVCI (Pan et al., 2022) and Kite (Li and Riva, 2018) can only recognize user intents limited to predefined UI trace collections. Voicify (Vu et al., 2023) is dependent on the application developer’s integration of deep links; applications with limited deep links exhibit lower feature coverage. When using search engines or AI tools like GPT for tutorial information, the results can be irrelevant, outdated (Venkatesh et al., 2022), fabricated (Shen et al., 2023), or low-quality (Burns et al., 2022).
The other focus is the understanding of mobile GUI semantics, which is crucial for widget identification and context comprehension. A wealth of research focuses on providing semantic text annotations for UI elements (Li et al., 2020b; He et al., 2021; Chen et al., 2022b; Wang et al., 2023). Other research efforts include generating full-page text descriptions (Wang et al., 2021), image-to-text conversion (Lee et al., 2022), and GUI embedding (Li et al., 2021). Translating GUI images and hierarchical information into natural language descriptions without information loss is vital for operation mapping, as it enables textual instructions to align with GUI operations.
To tackle these multifaceted knowledge demands, our system employs specialized agents, each focusing on a unique task: the Analysis Agent for understanding user intent from various text inputs through supplementing context information, the Retrieval Agent for managing online tutorial data, and the Mobile Semantics Agent for extracting UI semantics. These agents collaboratively build an open, accumulative knowledge base, enabling effective task automation on smartphones.
@@ 2.3. LLM-Driven Multi-Agent Frameworks
Large Language Models (LLMs) have been widely applied in decision-making tasks recently. For one thing, some studies focus on the reasoning capabilities of LLMs. ReAct (Yao et al., 2022) investigates the generation of reasoning traces and actions, and ExpeL (Zhao et al., 2023) delves into the realm of autonomous experiential learning. For another, increasing works aimed at enhancing the LLMs’ interactive capabilities with external environments. Toolformer (Schick et al., 2023) enables LLMs to interact with APIs, whereas Palo et al. (Palo et al., 2023) explores their role within reinforcement learning agents. Additionally, many works extend the utility of LLMs across different modalities. “Do As I Can, Not As I Say” (Ahn et al., 2022) introduces a grounding mechanism, ProgPrompt (Singh et al., 2023) targets robotic task planning, and VisualGPT (Chen et al., 2022a) examines multi-modal adaptability.
Moreover, several works address the issue of multi-agent cooperation driven by LLMs. MetaGPT (Hong et al., 2023) addresses illusion issues in complex tasks. Other frameworks, such as those proposed by Du et al. (Du et al., 2023), Liang et al. (Liang et al., 2023), Wei et al. (Wei et al., 2023) and Zhang et al. (Zhang et al., 2023), explore various aspects ranging from collective decision-making to divergent thinking.
Prior work has shown the value of multi-agent frameworks in solving complex tasks. In this paper, the workflow of PromptRPA is divided into several sub-tasks, each assigned to a specialized agent. By optimizing their performance individually and fostering collaboration among them, these agents collectively enhance the overall task execution.
2.4. Optimizing the System Based on User Feedback
@@
Interactive machine learning (IML) (Fails and Olsen, 2003; Dudley and Kristensson, 2018; Amershi et al., 2014) aims to achieve complementary benefits by integrating human perception and intelligence with the computational power of computers (Deng et al., 2020). The framework has seen extensive application in fields requiring annotation (Sperrle et al., 2019), classification (Berg et al., 2019; de Medeiros et al., 2020), and intricate data analysis (Carney et al., 2020).
In line with the principles of IML, PromptRPA employs an ensemble of agents that evolve through user feedback. To minimize the user’s cognitive load, PromptRPA supports feedback ranging from chats to simple text edits and hands-on demonstrations.
3. From Textual Prompts to RPA
Users approach tasks with a spectrum of understanding. Some may only have a conceptual need but lack the specifics on how to accomplish it, while others might already be armed with key procedural steps and are looking to automate them. In either case, there exists a gap between what users know or wish to achieve and the realization of these tasks through automation. Textual prompts serve as a vital bridge, which can express user intent and be transformed into structured sequences of operations, thereby enabling task automation.
In this section, we explore how textual prompts transition into RPA tasks. We introduce a formalized RPA task model to better understand this process. The model outlines the entire lifecycle of an RPA task, from its start to its completion, emphasizing the role of textual prompts in initiating and guiding the automation process. We then analyze the challenges within the process and outline the design principles that underpin our system’s architecture.
3.1. RPA Task Model
Robotic Process Automation (RPA) has historically been viewed as a tool designed to perform automated tasks by simulating human interactions with computer systems (Van der Aalst et al., 2018). This traditional definition focused largely on mimicking pre-defined, rule-based actions on graphical user interfaces.

### v2

# Prompt2Task: Automating UI Tasks on Smartphones from Textual Prompts

@@ Abstract.
UI task automation enables efficient task execution by simulating human interactions with graphical user interfaces (GUIs), without modifying the existing application code. However, its broader adoption is constrained by the need for expertise in both scripting languages and workflow design. To address this challenge, we present Prompt2Task, a system designed to comprehend various task-related textual prompts (e.g., goals, procedures), thereby generating and performing the corresponding automation tasks. Prompt2Task incorporates a suite of intelligent agents that mimic human cognitive functions, specializing in interpreting user intent, managing external information for task generation, and executing operations on smartphones. The agents can learn from user feedback and continuously improve their performance based on the accumulated knowledge. Experimental results indicated a performance jump from a 22.28% success rate in the baseline to 95.24% with Prompt2Task, requiring an average of 0.69 user interventions for each new task. Prompt2Task presents promising applications in fields such as tutorial creation, smart assistance, and customer service.
UI task automation, UI navigation, natural language, UI understanding, large language models
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Human-centered computing Graphical user interfacesccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
UI task automation has emerged as a pivotal technology, significantly boosting productivity and operational efficiency (Li et al., 2017). Its ability to simulate human interactions on the graphical user interface (GUI) allows for broad compatibility with third-party applications, without necessitating modifications to existing code (Deka et al., 2017; Jovanović et al., 2018). However, task automation primarily relies on predefined operation sequences (Ribeiro et al., 2021; Herm et al., 2023). Customizing these sequences often requires specialized skills in scripting languages and workflow design, which is a significant barrier to adoption, particularly for users without a technical background (Aguirre and Rodriguez, 2017). Additionally, the fixed operation sequences have difficulty adapting to the frequently updated GUI, making them susceptible to becoming quickly outdated.
Figure 1. The workflow of Prompt2Task. In the first stage, users input task-related textual prompts, and the agent collects relevant textual information to plan the task. In the second stage, the agent converts the text into formally defined instructions. In the third stage, the agent maps these instructions to operations, ultimately executing the automation task on the smartphone. Users can intervene in the agents’ decisions, especially when help is requested. Blue arrows guide the process from prompts to task automation, while green arrows indicate the knowledge accumulation through user interactions, thereby enhancing the future performance of the agents.
@@
To address this gap, we introduce Prompt2Task, which employs a text-prompt-driven task automation approach implemented on smartphones. Prompt2Task is capable of accepting unrestricted textual descriptions for UI tasks—from concrete operational instructions like “open settings, locate WLAN, and…” to high-level functional requests such as “change the ringtone”. Prompt2Task then automatically translates these textual prompts into operation sequences on GUI, freeing users from the complexity of designing automation scripts themselves.
@@
Prompt2Task serves as an intelligent navigation system on smartphones. It interprets user intent from textual descriptions and matches them to its internal “map” for route planning and decision-making, thereby automatically generating task flows. Throughout the process, Prompt2Task faces several challenges. Firstly, it needs to accommodate diverse and unclear textual inputs, requiring Prompt2Task to infer intentions from vague or incomplete instructions (Wen et al., 2001). Secondly, the “map” may be incomplete, especially lacking detailed route information for less common tasks or newly added application functions (Islam et al., 2010). This issue is difficult to address even with state-of-the-art Large Language Models (LLMs), which may generate incorrect or misleading information in the GUI domain (Ji et al., 2023; Wang et al., 2023). Thirdly, it is challenging to adapt to frequent updates in interfaces and applications; as they evolve, the routes Prompt2Task relies on may also require updates (Venkatesh et al., 2022). These challenges span across different stages of the process, requiring the application of various types of intelligence and knowledge to address them effectively.
@@
To meet these challenges, we employ a multi-agent framework (Panait and Luke, 2005) for Prompt2Task. The framework allocates specialized agents to sub-tasks within different stages of the process, with each agent emulating specific human cognitive functions. As illustrated in Figure 1, Prompt2Task spans three stages—from information collection and instruction generation to operation mapping. Specifically, information collection involves gathering all necessary descriptions of functions and steps for the automation task. Instruction generation entails creating formalized instructions to guide operations, and operation mapping refers to translating text-based instructions into operations on smartphones to ultimately complete the task.
@@
Two core elements are critical for Prompt2Task’s performance: a comprehensive knowledge base (Venkatesh et al., 2022), which is our smartphone “map”, and stable decision-making (Zhong et al., 2021), which ensures the agents’ efficient and collaborative task execution. Therefore, we create an evolving knowledge base that expands as it accumulates data and user interactions. We introduce three key agents to meet the specific knowledge demands of UI task automation: the Analysis Agent, which identifies user intents and actively fills in missing contextual parameters from user inputs; the Retrieval Agent, which sources online tutorials to guide task automation; and the Grounding Agent, which conducts in-depth semantic analysis of mobile GUI for better alignment with textual inputs and supports unplanned operations upon updates or errors. To enhance decision-making, the Assessment Agent scrutinizes decisions made by other agents, serving as an internal audit mechanism to improve accuracy and reliability. Beyond internal audit, the agents are seamlessly integrated into a user interface, allowing for user oversee and intervention to enhance decisions. User intervention is optional and involves chatting, selecting, editing, or demonstrating to modify and confirm the agents’ decisions. User usage and intervention create a feedback loop for continuous knowledge accumulation and system improvement, thereby optimizing efficiency for future tasks.
@@
To evaluate Prompt2Task, we conducted a performance evaluation with 2,500 textual prompts crowd-sourced. The results showed that Prompt2Task increased the task success rate from a baseline of 22.28% to 95.24%, requiring only an average of 0.69 user interventions for each new task. Moreover, it was proved that the accumulation of user usage and intervention improved system performance. Furthermore, a user study involving both skilled and unskilled smartphone users confirmed Prompt2Task’s effectiveness in assistance and its low interaction cost. Due to the low-cost translation from text to automation tasks, Prompt2Task has promising applications in generating interactive tutorials, enhancing smart assistance, and improving online customer service.
Specifically, the contributions of this work are three-fold.
  • We propose Prompt2Task, a multi-agent system that accepts unrestricted textual expressions of task requirements and automates corresponding UI tasks. To improve the success rate, Prompt2Task incorporates several innovative strategies: retrieving online tutorials, integrating necessary user interventions, and continuously improving agent capabilities through the accumulation of knowledge from use.
  • We propose an effective approach to minimize the cost of user interventions. Prompt2Task presents results stage-by-stage, allowing users to intuitively interact with the system through dialogues, selections, edits, and demonstrations. A dedicated agent employs GPT-4 to evaluate the confidence of predicted operations, requesting user assistance only when necessary.
  • We create a Chinese dataset  [^1] The dataset is at https://github.com/PromptRPA/Prompt2TaskDataset. that covers a wide range of UI tasks and collects diverse user descriptions of the tasks. The dataset includes 2,500 textual prompts, spanning 100 tasks across 10 application domains, with a total of 533 instructions and 543 operations.
@@ 2. Related Work
Our work intersects with four research areas: prior UI task automation systems, the knowledge required for UI task automation, LLM-driven multi-agent frameworks, and incorporating user feedback for agent improvement.
2.1. Prior UI Task Automation Systems
Prior UI task automation systems on smartphones typically work by replaying predefined interaction flows on graphical user interfaces, thereby completing the automation tasks. Early systems like SUGILITE (Li et al., 2017) and APPINITE (Li et al., 2018) utilized programming by demonstration to enable users to create interaction flows. However, for complex or unfamiliar functions, creating flows is challenging, especially for beginners and the elderly, who may lack essential task execution skills (Chen et al., 2017; Jin et al., 2022). Subsequent developments introduce interactive, context-aware tutorials for GUI applications (Fraser et al., 2019; Pongnumkul et al., 2011; Li et al., 2015; Zhong et al., 2021; Li, 2021). However, these systems primarily rely on pre-configured flows by developers (Wang et al., 2014). When accommodating new tasks or when interface updates cause the original flows outdated (Wu et al., 2023), these systems require users to have specialized skills in script design (Wewerka et al., 2020).
Recent efforts have focused on translating natural language into operation sequences on smartphones. Notable contributions include Branavan et al. (Branavan et al., 2009) and Li et al. (Li et al., 2020a). However, they require detailed instructions for each step, necessitating that users have a thorough understanding of task execution. While there are systems that support general task requests, limitations exist, such as UGIF’s reliance on a pre-collected tutorial dataset (Venkatesh et al., 2022) and Voicify’s dependence on app developers’ implementation of deep links (Vu et al., 2023), failing to meet a wide range of user needs. Additionally, they often function as “black boxes”, which can autonomously complete tasks when functioning correctly but may also produce uncontrollable and erroneous results (Venkatesh et al., 2022).
2.2. Knowledge Required for UI Task Automation
@@
Addressing user input presents challenges in accuracy and scope. User inputs are often vague (Jansen et al., 1998), requiring supplementary contextual information for accurate recognition of user intent (Grudin and Jacques, 2019). In terms of scope, existing UI trace datasets such as Rico (Deka et al., 2017), UGIF-DataSet (Venkatesh et al., 2022), and PixelHelp (Li et al., 2020a) are pre-collected and they are insufficient to meet a wide array of dynamic user needs (Beneteau et al., 2019). AutoVCI (Pan et al., 2022) and Kite (Li and Riva, 2018) can only recognize user intents limited to predefined UI trace collections. When using search engines or AI tools like GPT for tutorial information, the results can be irrelevant, outdated (Venkatesh et al., 2022), fabricated (Shen et al., 2023), or low-quality (Burns et al., 2022).
The other focus is the understanding of mobile GUI semantics, which is crucial for widget identification and context comprehension. A wealth of research focuses on providing semantic text annotations for UI elements (Li et al., 2020b; He et al., 2021; Chen et al., 2022b; Wang et al., 2023). Other research efforts include generating full-page text descriptions (Wang et al., 2021), image-to-HTML conversion (Lee et al., 2022), and GUI embedding (Bai et al., 2021; Li et al., 2021). Translating GUI images and hierarchical information into natural language descriptions while preserving key task-relevant information is crucial for operation mapping, as it enables textual instructions to align effectively with GUI operations.
@@ 2.3. LLM-Driven Multi-Agent Frameworks
Large Language Models (LLMs) have been widely applied in decision-making tasks recently. For one thing, some studies focus on the reasoning capabilities of LLMs. ReAct (Yao et al., 2022) investigates the generation of reasoning traces and actions, and ExpeL (Zhao et al., 2023) delves into the realm of autonomous experiential learning. For another, increasing works aimed at enhancing the LLMs’ interactive capabilities with external environments. Toolformer (Schick et al., 2023) enables LLMs to interact with APIs, whereas Palo et al. (Palo et al., 2023) explores their role within reinforcement learning agents. “Do As I Can, Not As I Say” (Ahn et al., 2022) introduces a grounding mechanism, ProgPrompt (Singh et al., 2023) targets robotic task planning, and VisualGPT (Chen et al., 2022a) examines multi-modal adaptability. Moreover, several works focus on multi-agent cooperation driven by LLMs. The frameworks proposed by Hong et al. (Hong et al., 2023b), Du et al. (Du et al., 2023), Liang et al. (Liang et al., 2023), Wei et al. (Wei et al., 2023) and Zhang et al. (Zhang et al., 2023a), explore various aspects ranging from collective decision-making to divergent thinking.
Prior work has shown the effectiveness of multi-agent frameworks in addressing complex tasks, which we have applied in designing Prompt2Task.
2.4. Incorporating User Feedback for Agent Improvement
@@
Interactive machine learning (IML) (Fails and Olsen, 2003; Dudley and Kristensson, 2018; Amershi et al., 2014) aims to achieve complementary benefits by integrating human perception and intelligence with the computational power of computers (Deng et al., 2020). The framework has seen extensive application in fields requiring annotation (Sperrle et al., 2019), classification (Berg et al., 2019; de Medeiros et al., 2020), and intricate data analysis (Carney et al., 2020). In line with the principles of IML, Prompt2Task employs an ensemble of agents that evolve through user feedback. The key problem is how to incorporate minimal, low-cost user intervention to achieve the maximum improvement in system performance.
3. Problem of UI Task Automation